Ex2: A beam with a weight of 500 N is connected to the wall with a hinge. The beam is horizontal and is held in place by a cable that is tied off to a wall. The angle between the cable and beam is 30 degrees. An object with a weight of 200 N is hanging at the end of the beam. a.What is the tension in the cable? b.What is the magnitude and direction of the force exerted on the beam by the hinge?

physics
Answer #1

a] Let the tension be T,

As it is at equilibrium, net torque about hinge = 0

Wbeam*L/2 + Wobject*L - T*L*sin theta = 0

500/2 + 200 - T* sin 30 degree = 0

T = [500/2 + 200]/sin 30 degree = 900 N answer